2.1. Let P(A) = 0.5, P(B) = 0.6 and P(A ∩ B) = 0.3. Find P(B¯), P(A ∩ B¯) and P(A ∪ B)


1
Expert's answer
2021-02-24T06:49:19-0500

"P(\overline{B})=1-P(B)=0.4\\
P(A\cap \overline{B})=P(A)-P(A\cap B)=0.2\\
P(A\cup B)=P(A)+P(B)-P(A\cap B)=0.8"